OpenStudy (anonymous):

Help Simplify, rationalize denominators: 6 ^3radical 3 -2 ^3radical 3

OpenStudy (anonymous):

\[6\sqrt[3]{3}- 2\sqrt[3]{3}\]

OpenStudy (anonymous):

\[4\sqrt[3]{3}\]

OpenStudy (anonymous):

thank you!